OpenStudy (anonymous):

factor the trinomial completely ........25x^3-55x^2-60x

OpenStudy (campbell_st):

take 5x out as a common factor 1st 5x(5x^2 - 11x - 12) next factor the trinomial in the brackets.

OpenStudy (anonymous):

@campbell_st, how would I do that there are no GCF for 5x^2-11x-12...I'm confused

OpenStudy (anonymous):

@campbell_st, how would I do that there are no GCF for 5x^2-11x-12...I'm confused

OpenStudy (campbell_st):

here is my method for a trinomial like this or any quadratic in the form ax^2 + bx + c multiply a and c 5 x -12 = -60 find the factors of -60 that add to - 11...

OpenStudy (campbell_st):

are you able to find the factors..?

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

10 and 1?

OpenStudy (campbell_st):

then it is  you'll find that there are common factors in the numerator that will cancel in the denominator.

OpenStudy (campbell_st):

10 x 1 = 10 find 2 numbers that multiply to -60 and add to -11 try 6 x -10 = -60 6 + -10 = -4 not them 3 x -20 = -60 3 + -20 = -17 not them 5 x - 12 = -60 5 + -12 = -7 not them any ideas..?

OpenStudy (anonymous):

oh multiply ok I missed that some kind of way

OpenStudy (anonymous):

ok... I got nothing... I cant think of anything that would be multiplied to get 60 and added to get 11

OpenStudy (campbell_st):

well you need something between 3 x -20 and 5 x - 12... any thoughts...?

OpenStudy (anonymous):

im sorry..I'm thinking about it

OpenStudy (anonymous):

sorry...I have no idea what I'm looking for here...I dont know what's in between 3x-20 and 5x-12.. i just dont get this

OpenStudy (campbell_st):

try 4 and - 15.... so you have  the solution for the trinomial is done... all you need to do is combine the solution.

OpenStudy (anonymous):

ok I kind of understand

OpenStudy (anonymous):

so first i look for GCF then divide it by that?

OpenStudy (anonymous):

what do u mean combine it

OpenStudy (campbell_st):

5x(x -3)(5x + 4)